the specific heat of ethanol is 2.46 j/gc. when 2555 j of heat are added to 32.8 g ethanol, a final temperature of 77.5 c is reached. what was the initial temperature of ethanol
Answer #1

q=mcΔt,

where q is heat in Joules, m is mass in g, c is specific heat capacity in J/g0C, and Δt is change in temperature: Δt=(Tfinal−Tinitial)

q=2555J

m=32.8 g

c=2.46 J/g⋅C∘

Tinitial=?

Tfinal=77.5 0C

Δt=77.5-Tinitial

2555J =32.8g*2.46J/g⋅C∘*(77.5-Tinitial)

77.5-Tinitial=2555J/(32.8g*2.46J/g⋅C∘) =31.60C

Tinitial=77.5-31.6=45.90C

THANK YOU